# Learning Encodings by Maximizing State Distinguishability: Variational Quantum Error Correction

### v1

@@ E.3 Extended Stability Analysis
(a) Depolarizing noise with noise strength $p=0.1$.
(b) Asymmetric depolarizing noise with noise strength $p=0.1$ and asymmetry $c=0.5$ following Eq. 41.
Figure 17: Stability of standard QEC and VarQEC codes with increasing noise strength for (a) symmetric and (b) asymmetric depolarizing noise. All data points represent the respective worst-case distinguishability loss $\overline{\mathcal{D}}$. The [[$5,1$]] VarQEC code is always the same static one trained on a noise strength $p=0.1$. The other [[$5,1$]] VarQEC code is dynamically trained on the respective denoted noise configuration.
(a) Depolarizing noise with noise strength $p=0.1$ on $n-1$ of the qubits, and with a noise strength as denoted on the $x$-axis on the remaining qubit. For the QEC codes, the deviating qubit is associated with the original data qubit, while the unencoded baseline employs the best available wire.
Figure 18: Stability of standard QEC and VarQEC codes in setups of varying (a) noise strength on different qubits and (b) noise channels. All data points represent the respective worst-case distinguishability loss $\overline{\mathcal{D}}$ and optionally average-case loss $\mathcal{D}$. The [[$5,1$]] VarQEC code is always the same static one trained on depolarizing noise with noise strength $p=0.1$ on all qubits. The other [[$5,1$]] VarQEC code is dynamically trained on the respective noise configuration.
@@
We now analyze the stability of the VarQEC codes and the adaptability of the VarQEC procedure. Therefore, for both symmetric and asymmetric noise, we vary the noise strength from $p=0.05$ to $p=0.25$. Similar to the main part, we compare the unencoded baseline, the [[$5,1,3$]] perfect code, a static VarQEC code trained on noise strength of $p=0.1$, and VarQEC codes dynamically trained on the respective noise strengths. For the depolarizing noise in Fig. 17(a), we observe that all QEC codes perform equally well up to the physical noise threshold of about $p=0.15$. After that, both the [[$5,1,3$]] code and static VarQEC code perform worse than just a physical qubit without encoding. The dynamically trained VarQEC code at least performs on par with the purely physical setup, but also can not improve upon it. For the asymmetric depolarizing noise in Fig. 17(b), the [[$5,1,3$]] code requires a physical threshold of again approximately $p=0.15$. However, both VarQEC codes can significantly improve upon this and still work for $p=0.25$. While dynamically tailoring the VarQEC code brings slight advantages for some configurations, the differences are not significant.
@@
We extend this analysis by evaluating the scenario of non-uniform noise strength between the involved qubits in Fig. 18(a), which is typically resembled in actual hardware due to manufacturing quality [77]. More concretely, we vary the strength of depolarizing noise on the qubit corresponding to the initial data wire from values $p=0.05$ to $p=0.2$. All other involved ancilla qubits exhibit a depolarizing noise of $p=0.1$. To compare to a harder baseline, for the case of a single physical qubit, we select the lower noise strength. As expected, the [[$5,1,3$]] perfect code outperforms the purely physical setup up until a threshold noise strength of approximately $0.175$. The static version of the [[$5,1$]] VarQEC code performs on par for all setups. The [[$5,1$]] VarQEC dynamically trained on the respective noise configurations does not drop beyond the baseline, even above the physical threshold.
@@
In Fig. 18(b) we perform some transferability analysis of the [[$5,1$]] VarQEC code that was trained on depolarizing noise of strength $p=0.1$. As evaluated in Section B.3, this code exhibits a potential code distance of $d=3$, and therefore also should apply to other noise channels. Indeed, the transferred code exhibits distinguishability losses that are significantly beyond the unencoded baselines (compare Fig. 14). Moreover, for amplitude damping and combined amplitude and phase damping noise, the results are again competitive with the [[$5,1,3$]] perfect code. We have to highlight that one can further improve upon this by tailoring the codes explicitly to the respective noise channels. However, these results still demonstrate that the learned VarQEC codes partially generalize to other noise models. This might allow for some warm-start parameter initialization approaches [65], indicating that models do not always need to be trained from scratch but rather only finetuned.
@@ E.4 Restricting Connectivity Structure and Fault Resilience
$a_{1}$$a_{2}$$a_{3}$$\ d$$a_{4}$
@@
Most quantum hardware platforms do not support all-to-all connectivity, but rather assume some restricted topology. In principle, it is possible to compile a quantum circuit to the respective layout, which, however, typically requires routing with e.g. swap gates. We can entirely avoid this for the VarQEC procedure by initially selecting a variational ansatz that is adapted to the hardware connectivity restrictions. Therefore, we now consider the four different layouts in Fig. 19, i.e.,  “dense”, “square”, “star”, and “hexagonal”. The randomized entangling ansätze are configured to inherently apply to these restrictions.
@@
[[$5,1,3$]]] Perfect $6$ $9$ $15$ $12$ $18$
[[$5,1$]] VarQEC $12$ $10$ $10$ $12$ $16$
Table 5: Number of two-qubit gates for [[$5,1,3$]] perfect code compiled to the different layouts in Fig. 19 and smallest number of two-qubit blocks for which a [[$5,1$]] VarQEC code achieves the optimal worst-case distinguishability loss of $\overline{\mathcal{D}}\approx 0.106$ for depolarizing noise with $p=0.1$.
For the results in Table 5, we trained [[$5,1$]] VarQEC codes with $6$ to $20$ two-qubit blocks on depolarizing noise. For each layout, we report the respective smallest code instance that produces the optimal distinguishability loss of $\overline{\mathcal{D}}\approx 0.106$. In addition to that, we compile the [[$5,1,3$]] perfect encoding circuit [103] for the respective layout, using the qiskit.transpile method [45], and report the respective number of two-qubit gates. This significantly increases the circuit depth, while for the VarQEC codes, the circuit complexity only needs to increase slightly with lower connectivity density.
@@
Table 6: Worst-case distinguishability loss $\overline{\mathcal{D}}$ for thermal relaxation noise with duration $t=10.0\mu\text{s}$ with $T_{1}=200\mu\text{s}$ and $T_{2}=100\mu\text{s}$ using an increasing number of two-qubit blocks in [[$5,1$]] VarQEC encoding circuits.
For realistically modeling idle qubits, one typically resorts to the thermal relaxation noise channel. Therefore, we repeat the above analysis on the VarQEC circuits for this setup with $T_{1}=200\mu\text{s}$, $T_{2}=100\mu\text{s}$, $t=10.0\mu\text{s}$, and report results in Table 6. The results underline that the VarQEC can easily adapt to sparse connectivity, with just a slight increase in circuit depth.
Figure 20: Resilience of standard QEC and VarQEC codes in setups of non-ideal encoding operations under symmetric depolarizing noise of strength $p=0.1$. All data points represent the respective worst-case distinguishability loss $\overline{\mathcal{D}}$. The solid lines represent static VarQEC codes trained under the assumption of ideal encoding operations, adapted to the respective layouts. The dashed lines correspond to the [[$5,1,3$]] QEC code, compiled for the respective connectivity structure. The unencoded baseline, which does not include any two-qubit operations, is at approximately $0.133$. The encoding circuit depths of the respective codes are summarized in Table 5.
Last but not least, we consider the noise-resilience of the encodings under the assumption of restricted connectivity Fig. 20. Similar to the main part, we evaluate this by adding correlated depolarizing noise of increasing strength after each two-qubit gate. We consider symmetric depolarizing noise, i.e. under the assumption of ideal encoding operations, the [[$5,1,3$]] code is optimal. For “full”, “dense”, and “’square’ connectivity layouts, the compiled codes are more resilient than the VarQEC codes, mainly due to the circuits being shallower. However, the VarQEC codes shine under sparse connectivity structures like especially “hexagonal”, where this trend reverses. We want to emphasize that, e.g. for asymmetric depolarizing noise, the VarQEC codes would exhibit superior noise-resilience on all layouts, as already under full-connectivity assumptions, there was quite some performance gap as analyzed in Fig. 6(b).
@@ F.1 General Experimental Setup
Figure 21: Distinguishability loss for different QEC codes on thermal relaxation noise modeling the median error properties for (a) an IBMQ device and (b) an IQM device. All bars visualize the average-case loss $\mathcal{D}$ and worst-case loss $\overline{\mathcal{D}}$. For the IBMQ noise model, we compare the unencoded baseline and VarQEC codes with $n=3$ to $n-5$ physical qubits. The [[$4,1$]] VarQEC code for this setup is plotted in Fig. 23. For the IQM noise model, we compare the unencoded baseline to the results of [[$4,1$]] VarQEC codes with different layout alignments. The codes are trained in classical simulation and afterwards deployed on quantum hardware, as denoted by the experiments in this section.
@@
We compared the hardware results to the training predictions on the distinguishability loss, as depicted in Fig. 21. While the hardware measurements reported in Figs. 22 and 24 exhibit quantitatively higher distinguishability loss than the simulation predictions, they still demonstrate beyond break-even performance as we discuss in the following, corroborating the observations from the main text in Section 6.
@@ F.2 Extended Results on IBMQ Decides
Figure 22: Reconstructed distinguishability loss for the $8$ patches of the experiment with the [[$4,1$]] VarQEC code on the ibm_marrakesh device. The qubits have been idled for $5.0\mu\text{s}$ for this experiment. For each patch, the first four bars indicate the unencoded baseline losses for the respective physical qubits, while the fifth one denotes the loss for the entire encoded patch. The losses were estimated with mitigated state tomography using $10000$ shots per circuit execution, and the error bars denote 5th and 95th percentile ranges estimated by cluster re-sampling. The alignment and layout of physical wires to logical patch is also summarized in Table 7, and the circuit realizing the encoding is given in Fig. 23.
logical patch $\ d$$a_{2}$$a_{1}$$a_{3}$$a_{4}$
@@ I II III IV V VI VII VIII
Table 7: Physical wire allocation of the encoding patches for the experiment on the IBMQ device ibm_marrakesh depicted in Fig. 22. The device uses a Heron r2 chip with 156 qubits, the physical wire indices refer to the numbering in the IBMQ documentation [79]. The connectivity is adapted to the hexagonal connectivity structure, each single 5-qubit encoding patch is allocated as depicted on the right. Hereby, $d$ denotes the initial physical data qubit to be protected, and $a_{1}$, $a_{2}$, $a_{3}$, and $a_{4}$ are ancilla qubits. For the experiments with encodings on 4 and 3 wires, the ancilla qubits $a_{4}$ and $a_{3}$ are dropped, respectively.
@@
Figure 23: Circuit ansatz implementing the [[$4,1$]] VarQEC code. The parameters were trained on the noise setup for the iqm_marrakesh experiment that is analyzed in Fig. 22. The circuit is an instance of a randomized entangling ansatz (RAG, see Section D.1) restricted to the “hexagonal” four-qubit layout with the basis gate set of the IBM devices. It contains four two-qubit blocks, with overall $36$ parameters. The uppermost wire corresponds to the initial data qubit.
@@
We conducted experiments on the ibm_marrakesh device [79], which features a $156$-qubit Heron r2 chip. The median properties of the qubits were identified as $T_{1}=180\mu\text{s}$ and $T_{1}=120\mu\text{s}$. Due to additional sources of error, such as noise introduced during the encoding operations and measurement delays, we trained the encoding with a noise duration of $t=10.0\mu\text{s}$ and deployed it on the hardware for $t=5.0\mu\text{s}$ (compare also Section 6).
@@
We trained and deployed VarQEC codes with $n=3$ to $n=5$ physical qubits. The experiments required approximately $12$ QPU minutes for the $3$-qubit code, $25$ QPU minutes for the $4$-qubit code, and $72$ QPU minutes for the $5$-qubit code, accessed via the IBM Quantum premium plan. The logical qubit allocations for the [[$3,1$]] and [[$5,1$] VarQEC codes are summarized in Table 7. The gates used in our ansätze were aligned with the native gate set of the device, employing CZ gates for two-qubit operations and parameterized single-qubit rotations implemented as $R_{zxz}(\theta,\phi,\xi)=R_{z}(\xi)R_{x}(\phi)R_{z}(\theta)$. The results, depicted in Fig. 8(b) of the main text, indicate a decrease in the distinguishability loss with an increasing number of physical qubits, thus demonstrating improved error correction performance with larger codes.
@@
Detailed results for the $8$ individual patches of the [[$4,1$]] VarQEC code are provided in Fig. 22. We note both worst-case and average-case distinguishability losses, with error bars indicating the 5th and 95th percentile ranges evaluated by cluster re-sampling. The experiments were conducted with $10000$ shots for each circuit in the mitigated state tomography process. The concrete ansatz for the [[$4,1$]] VarQEC code, consisting of four two-qubit blocks, is illustrated in Fig. 23. We observed that while the VarQEC code does not outperform the best physical qubit in every individual patch, especially in patches with significant qubit quality fluctuations, the overall results indicate a reduction in the distinguishability loss on average. This is particularly pronounced in patches where the qubit quality is more uniform (e.g., patch IV). In some patches, the logical results did not outperform the best physical qubit, which we attribute to variances in qubit quality and, in one case, a defective two-qubit connection (patch VIII).

### v2

@@ E.3 Extended Stability Analysis
We now analyze the stability of the VarQEC codes and the adaptability of the VarQEC procedure. Therefore, for both symmetric and asymmetric noise, we vary the noise strength from $p=0.05$ to $p=0.25$. Similar to the main part, we compare the unencoded baseline, the $[[5,1,3]]$ perfect code, a static VarQEC code trained on noise strength of $p=0.1$, and VarQEC codes dynamically trained on the respective noise strengths. For the depolarizing noise in figure˜17(a), we observe that all QEC codes perform equally well up to the physical noise threshold of about $p=0.15$. After that, both the $[[5,1,3]]$ code and static VarQEC code perform worse than just a physical qubit without encoding. The dynamically trained VarQEC code at least performs on par with the purely physical setup, but also can not improve upon it. For the asymmetric depolarizing noise in figure˜17(b), the $[[5,1,3]]$ code requires a physical threshold of again approximately $p=0.15$. However, both VarQEC codes can significantly improve upon this and still work for $p=0.25$. While dynamically tailoring the VarQEC code brings slight advantages for some configurations, the differences are not significant.
@@
We extend this analysis by evaluating the scenario of non-uniform noise strength between the involved qubits in figure˜18(a), which is typically resembled in actual hardware due to manufacturing quality [83]. More concretely, we vary the strength of depolarizing noise on the qubit corresponding to the initial data wire from values $p=0.05$ to $p=0.2$. All other involved ancilla qubits exhibit a depolarizing noise of $p=0.1$. To compare to a harder baseline, for the case of a single physical qubit, we select the lower noise strength. As expected, the $[[5,1,3]]$ perfect code outperforms the purely physical setup up until a threshold noise strength of approximately $0.175$. The static version of the $((5,2))$ VarQEC code performs on par for all setups. The $((5,2))$ VarQEC dynamically trained on the respective noise configurations does not drop beyond the baseline, even above the physical threshold.
@@
In figure˜18(b) we perform some transferability analysis of the $((5,2))$ VarQEC code that was trained on depolarizing noise of strength $p=0.1$. As evaluated in section˜B.3, this code exhibits a potential code distance of $d=3$, and therefore also should apply to other noise channels. Indeed, the transferred code exhibits distinguishability losses that are significantly beyond the unencoded baselines (compare figure˜14). Moreover, for amplitude damping and combined amplitude and phase damping noise, the results are again competitive with the $[[5,1,3]]$ perfect code. We have to highlight that one can further improve upon this by tailoring the codes explicitly to the respective noise channels. However, these results still demonstrate that the learned VarQEC codes partially generalize to other noise models. This might allow for some warm-start parameter initialization approaches [69], indicating that models do not always need to be trained from scratch but rather only finetuned.
@@ E.4 Restricting Connectivity Structure and Fault Resilience
Most quantum hardware platforms do not support all-to-all connectivity, but rather assume some restricted topology. In principle, it is possible to compile a quantum circuit to the respective layout, which, however, typically requires routing with e.g. swap gates. We can entirely avoid this for the VarQEC procedure by initially selecting a variational ansatz that is adapted to the hardware connectivity restrictions. Therefore, we now consider the four different layouts in figure˜19, i.e., “dense”, “square”, “star”, and “hexagonal”. The randomized entangling ansätze are configured to inherently apply to these restrictions.
$a_{1}$$a_{2}$$a_{3}$d\ d\ $a_{4}$
@@
$[[5,1,3]]$ Perfect 66 99 $15$ $12$ $18$
$((5,2))$ VarQEC $12$ $10$ $10$ $12$ $16$
Table 5: Number of two-qubit gates for $[[5,1,3]]$ perfect code compiled to the different layouts in figure˜19 and smallest number of two-qubit blocks for which a $((5,2))$ VarQEC code achieves the optimal worst-case distinguishability loss of $\overline{\mathcal{D}}\approx 0.106$ for depolarizing noise with $p=0.1$.
For the results in table˜5, we trained $((5,2))$ VarQEC codes with $6$ to $20$ two-qubit blocks on depolarizing noise. For each layout, we report the respective smallest code instance that produces the optimal distinguishability loss of $\overline{\mathcal{D}}\approx 0.106$. In addition to that, we compile the $[[5,1,3]]$ perfect encoding circuit [111] for the respective layout, using the qiskit.transpile method [49], and report the respective number of two-qubit gates. This significantly increases the circuit depth, while for the VarQEC codes, the circuit complexity only needs to increase slightly with lower connectivity density.
@@
Table 6: Worst-case distinguishability loss $\overline{\mathcal{D}}$ for thermal relaxation noise with duration $t=10.0\mu\text{s}$ with $T_{1}=200\mu\text{s}$ and $T_{2}=100\mu\text{s}$ using an increasing number of two-qubit blocks in $((5,2))$ VarQEC encoding circuits.
Figure 20: Resilience of standard QEC and VarQEC codes in setups of non-ideal encoding operations under symmetric depolarizing noise of strength $p=0.1$. All data points represent the respective worst-case distinguishability loss $\overline{\mathcal{D}}$. The solid lines represent static VarQEC codes trained under the assumption of ideal encoding operations, adapted to the respective layouts. The dashed lines correspond to the $[[5,1,3]]$ QEC code, compiled for the respective connectivity structure. The unencoded baseline, which does not include any two-qubit operations, is at approximately $0.133$. The encoding circuit depths of the respective codes are summarized in table˜5.
For realistically modeling idle qubits, one typically resorts to the thermal relaxation noise channel. Therefore, we repeat the above analysis on the VarQEC circuits for this setup with $T_{1}=200\mu\text{s}$, $T_{2}=100\mu\text{s}$, $t=10.0\mu\text{s}$, and report results in table˜6. The results underline that the VarQEC can easily adapt to sparse connectivity, with just a slight increase in circuit depth.
Last but not least, we consider the noise-resilience of the encodings under the assumption of restricted connectivity figure˜20. Similar to the main part, we evaluate this by adding correlated depolarizing noise of increasing strength after each two-qubit gate. We consider symmetric depolarizing noise, i.e. under the assumption of ideal encoding operations, the $[[5,1,3]]$ code is optimal. For “full”, “dense”, and “’square’ connectivity layouts, the compiled codes are more resilient than the VarQEC codes, mainly due to the circuits being shallower. However, the VarQEC codes shine under sparse connectivity structures like especially “hexagonal”, where this trend reverses. We want to emphasize that, e.g. for asymmetric depolarizing noise, the VarQEC codes would exhibit superior noise-resilience on all layouts, as already under full-connectivity assumptions, there was quite some performance gap as analyzed in figure˜6(b).
@@ F.1 General Experimental Setup
Figure 21: Distinguishability loss for different QEC codes on thermal relaxation noise modeling the median error properties for (a) an IBMQ device and (b) an IQM device. All bars visualize the average-case loss $\mathcal{D}$ and worst-case loss $\overline{\mathcal{D}}$. For the IBMQ noise model, we compare the unencoded baseline and VarQEC codes with $n=3$ to n5n-5 physical qubits. The $((4,2))$ VarQEC code for this setup is plotted in figure˜23. For the IQM noise model, we compare the unencoded baseline to the results of $((4,2))$ VarQEC codes with different layout alignments. The codes are trained in classical simulation and afterwards deployed on quantum hardware, as denoted by the experiments in this section.
@@
We compared the hardware results to the training predictions on the distinguishability loss, as depicted in figure˜21. While the hardware measurements reported in figures˜22 and 24 exhibit quantitatively higher distinguishability loss than the simulation predictions, they still demonstrate beyond break-even performance as we discuss in the following, corroborating the observations from the main text in section˜6.
@@ F.2 Extended Results on IBMQ Decides
Figure 22: Reconstructed distinguishability loss for the 88 patches of the experiment with the $((4,2))$ VarQEC code on the ibm_marrakesh device. The qubits have been idled for $5.0\mu\text{s}$ for this experiment. For each patch, the first four bars indicate the unencoded baseline losses for the respective physical qubits, while the fifth one denotes the loss for the entire encoded patch. The losses were estimated with mitigated state tomography using $10000$ shots per circuit execution, and the error bars denote 5th and 95th percentile ranges estimated by cluster re-sampling. The alignment and layout of physical wires to logical patch is also summarized in table˜7, and the circuit realizing the encoding is given in figure˜23.
logical patch d\ d\ $a_{2}$$a_{1}$$a_{3}$$a_{4}$
@@ I II III IV V VI VII VIII
Table 7: Physical wire allocation of the encoding patches for the experiment on the IBMQ device ibm_marrakesh depicted in figure˜22. The device uses a Heron r2 chip with 156 qubits, the physical wire indices refer to the numbering in the IBMQ documentation [85]. The connectivity is adapted to the hexagonal connectivity structure, each single 5-qubit encoding patch is allocated as depicted on the right. Hereby, $d$ denotes the initial physical data qubit to be protected, and $a_{1}$, $a_{2}$, $a_{3}$, and $a_{4}$ are ancilla qubits. For the experiments with encodings on 4 and 3 wires, the ancilla qubits $a_{4}$ and $a_{3}$ are dropped, respectively.
@@
Figure 23: Circuit ansatz implementing the $((4,2))$ VarQEC code. The parameters were trained on the noise setup for the iqm_marrakesh experiment that is analyzed in figure˜22. The circuit is an instance of a randomized entangling ansatz (RAG, see section˜D.1) restricted to the “hexagonal” four-qubit layout with the basis gate set of the IBM devices. It contains four two-qubit blocks, with overall $36$ parameters. The uppermost wire corresponds to the initial data qubit.
@@
We conducted experiments on the ibm_marrakesh device [85], which features a $156$-qubit Heron r2 chip. The median properties of the qubits were identified as $T_{1}=180\mu\text{s}$ and $T_{1}=120\mu\text{s}$. Due to additional sources of error, such as noise introduced during the encoding operations and measurement delays, we trained the encoding with a noise duration of $t=10.0\mu\text{s}$ and deployed it on the hardware for $t=5.0\mu\text{s}$ (compare also section˜6).
@@
We trained and deployed VarQEC codes with $n=3$ to $n=5$ physical qubits. The experiments required approximately $12$ QPU minutes for the $3$-qubit code, $25$ QPU minutes for the $4$-qubit code, and $72$ QPU minutes for the $5$-qubit code, accessed via the IBM Quantum premium plan. The logical qubit allocations for the $((3,2))$, $((4,2))$ and $((5,2))$ VarQEC codes are summarized in table˜7. The gates used in our ansätze were aligned with the native gate set of the device, employing CZ gates for two-qubit operations and parameterized single-qubit rotations implemented as $R_{zxz}(\theta,\phi,\xi)=R_{z}(\xi)R_{x}(\phi)R_{z}(\theta)$. The results, depicted in figure˜8(b) of the main text, indicate a decrease in the distinguishability loss with an increasing number of physical qubits, thus demonstrating improved error correction performance with larger codes.
@@
Detailed results for the $8$ individual patches of the $((4,2))$ VarQEC code are provided in figure˜22. We note both worst-case and average-case distinguishability losses, with error bars indicating the 5th and 95th percentile ranges evaluated by cluster re-sampling. The experiments were conducted with $10000$ shots for each circuit in the mitigated state tomography process. The concrete ansatz for the $((4,2))$ VarQEC code, consisting of four two-qubit blocks, is illustrated in figure˜23. We observed that while the VarQEC code does not outperform the best physical qubit in every individual patch, especially in patches with significant qubit quality fluctuations, the overall results indicate a reduction in the distinguishability loss on average. This is particularly pronounced in patches where the qubit quality is more uniform (e.g., patch IV). In some patches, the logical results did not outperform the best physical qubit, which we attribute to variances in qubit quality and, in one case, a defective two-qubit connection (patch VIII).